OpenStudy (anonymous):

Heya! Double checking my answer since I think I am doing something wrong. Taking the positive infinite limit of (x^2)/sqr((x^4)+1). I divide by x^4 and eventually get 0/0 or DNE. When I check on the calulator however, it seems to be approaching 1. What am I doing wrong? :)

OpenStudy (anonymous):

The problem

OpenStudy (anonymous):

have you covered l'hopital's rule?

OpenStudy (anonymous):

I read about it online, but not explicitly, no

OpenStudy (anonymous):

We aren't "supposed" to know how to take derivativesin this section

OpenStudy (anonymous):

thats why I asked :P \[\lim_{x \rightarrow \infty} \frac{ x^2 }{ \sqrt{x+1} }\]

OpenStudy (anonymous):

oops to the power of 4 inside let me rewrite that

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}\frac{ x^2 }{ \sqrt{x^4+1} }\]

OpenStudy (anonymous):

lets think about this

OpenStudy (anonymous):

as we approach infinity do we care about the +1?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

so and this is not the way to write it out in your book but we can think of it as just the square root of x^4 on the bottom

OpenStudy (anonymous):

what is the square root of x^4?

OpenStudy (anonymous):

infinity in this case

OpenStudy (anonymous):

oh nvm

OpenStudy (anonymous):

in general not in this case

OpenStudy (anonymous):

is it (x^4)^(1/2)= x^(2) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can I do that? I thought because of the constant I can't apply the exponent to simplify

OpenStudy (anonymous):

technically we can't do it that way but I am showing you how I think of it. Now to do it more formally we divide both the numerator and the denominator by x^2 not x^4

OpenStudy (anonymous):

aha!

OpenStudy (anonymous):

remember that intially the divide by x^2 is outside of the square root so when you bring it in it changes to x^4

OpenStudy (anonymous):

ok. So a general question then for infinite limits. When we have a square root in the denominator and I am looking for the highest exponent to divide by, I should take sqrt(x^n) ?

OpenStudy (kainui):

I think you should try to think of this in an intuitive kind of way and think, hey, x^2 is only slightly smaller than sqrt(x^4+1) since if that +1 wasn't in there, it'd just be x^2 on bottom too! So as x gets larger and larger, that little 1 is going to be less noticeable until eventually, at infinity, the two values become pretty much the same. And anything divided by itself is just 1.

OpenStudy (anonymous):

I would lean more towards the thinking I walked you through and @Kainui just described again. It will help you understand it more than just a rinse and repeat procedure

OpenStudy (anonymous):

hmm. I can kind of visualise this in my head. but

OpenStudy (anonymous):

I'm still not exactly clear on where the x^2 divisor comes from

OpenStudy (anonymous):

whether it's a function of the particular ratio of this function or what

OpenStudy (anonymous):

like I said I took x^4 before, but that's obviously wrong

OpenStudy (anonymous):

well remember when we said what the square root of x^4 is?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so what I said was right, but I should also visualize

OpenStudy (anonymous):

thats where it came from looking at the new limit of \[\lim_{x \rightarrow \infty} \frac{ x^2 }{ x^2 }\]

OpenStudy (anonymous):

aha I see

OpenStudy (anonymous):

where I just subbed in what we said we cared about in the denominator

OpenStudy (anonymous):

its kinda weird how the way everyone thinks of it...is not the method you put down on paper lol

OpenStudy (anonymous):

but in terms of the steps I show..

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

so here's how I see it now

OpenStudy (anonymous):

first off, drop the 1

OpenStudy (anonymous):

then take sqrt(x^4) to get x^2

OpenStudy (anonymous):

x^2 / x^2 over x^2 / x^2

OpenStudy (anonymous):

which makes 1/1 =1

OpenStudy (anonymous):

am I seeing this correctly?

OpenStudy (anonymous):

yup thats how I think of it

OpenStudy (anonymous):

sweetness

OpenStudy (anonymous):

Alright well don't forget to close the question!

OpenStudy (anonymous):

yep, thanks